OpenStudy (dmezzullo):

Write an expression for the following real-world situation. Katie is moving to a new house. Small boxes hold 15 pounds each and medium boxes hold 25 pounds each. Write an expression for the total weight in the moving van if Katie uses "s" small boxes and "m" medium boxes. Part 2: Write a real-world situation of your own and a VARIABLE expression to go with it.

OpenStudy (anonymous):

Small boxes hold 15 lbs each Medium boxes hold 25 lbs each Small boxes = s Medium boxes = m 15 * s = 15s 25 * m = 25m 15s + 25m = total weight.

OpenStudy (anonymous):

Part 2: Write a real-world situation of your own and a VARIABLE expression to go with it. You can use a "per pound" situation like above. ex: Amy went to the store. Pears are 5 dollars per pound. Apples are 10 dollars per pound. Write an expression to solve the total weight. p = pears and a = apples. 5p + 10a = total pounds. ^Just use something else ;)

OpenStudy (dmezzullo):

ya

OpenStudy (swissgirl):

hmmm u didnt like may answer -_-